# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 5.1.1 Boltzmann Distribution
Human choice models as in (4) have been studied for a long time in various scientific fields such as psychology [249], economics [251] or behavioral science [47]. Accordingly, there are many different choice models to resort to for (5), which, in some cases, are the same models, just under different names. A popular class of such human choice models assumes every choice option $c$ to be equipped with a (latent) utility $u_{c}$, which the human perceives in a perturbed way. This perturbation is modeled by means of perturbation random variables $\epsilon_{c}$ that perturb the utility in an additive way, so that (4) becomes
@@
and we shall now stick to the RLHF translation from now on. These probabilities depend on the specific distributional assumptions that are made on the perturbation variables that only for specific cases lead to a closed-form of the right-hand sides of the latter displays. When stipulating a standard Gumbel distribution for the perturbations, one always obtains a closed form that is proportional to the exponential utility of the provided label:
@@
This is known as the Boltzmann distribution that also appears in a perhaps slightly modified version in various different subfields of machine learning and statistics. When restricting to discrete (choice) sets for $m(q)$, this distribution is also known as the multinomial logit model [251] or Gibbs measure [90], and as the Bradley-Terry model [38] when the choice sets consist of pairs. All of these also have a link to the Plackett-Luce model [173, 205], which is a probability distribution on the space of total orders or rankings (see [4] for details).
@@
This model is often used for various reasons. A particularly compelling reason is the closed analytic form, which in turn makes it possible to obtain a closed form for the gradient with respect to the utilities. Another reason is that this model satisfies Luce’s axiom of choice [205], which requires the probability of choosing an option from a pool of choice options not being affected by the presence or absence of other options in the pool. In this way, coherent decision-making is ensured, which, however, might be challenged as humans are likely not making fully rational decisions (see Section 4.2.1). [125] show that the usage of the Boltzmann distribution is justified by the principle of maximum entropy. More precisely, they show that it is the maximum entropy distribution over choices for a so-called satisficing human decision maker, i.e., one who is making in expectation a choice with an optimal reward up to some slack $\epsilon>0$.
@@
To build the bridge between reward learning and the modeling of human feedback, the Boltzmann distribution can be used by assuming that the utilities can be represented as a function of the reward function, usually featuring the return of a trajectory. More specifically, one assumes a grounding function $\psi$ that maps choice options (or labels) to the set of distributions over trajectories and sets the utility of a label $l$ as
$u_{l}:=\mathbb{E}_{\tau\sim\psi(l)}[R(\tau)]\,\text{.}$ (9)
@@
Table 1 in [125] provides an overview of the different grounding functions that lead to a specific feedback type. It is worth noting that one can also easily find a grounding function for the feedback type of a (partial) order over trajectories as considered, for instance, by [188]. Moreover, one can generalize this modeling approach by using (partial) segments instead of trajectories.
@@
Although this general human feedback model has been much in use and shown to be useful for the sake of human alignment, it is not without its critics (see [167] or Section 3.2.1 in [45]). This has led to different adaptions of the general model based on the Boltzmann distribution that will be discussed in the following. Moreover, we will also concisely review other human feedback models that have been in use besides the Boltzmann distribution, discuss relevant work on the consequences or robustness of human feedback model misspecification, and highlight contributions on varying the standard assumptions on the nature of the human feedback.
@@ 5.1.2 Human-Specific Rationality Coefficient
The higher $\beta$, the more (10) resembles a pointmass distribution modeling a highly rational human labeler (decision-maker) that is always able to identify the option with highest utility, while the lower $\beta$, the more (10) resembles a uniform distribution modeling a highly irrational human labeler (decision-maker) acting purely at random. Without this extension, the commonly considered Boltzmann distribution (or Bradley-Terry model in the common case of pairwise comparisons) in (8) assumes a rationality coefficient of $1$. [91] show in their experiments that the estimation of this coefficient can indeed positively influence reward learning. For the estimation, however, a calibration reward function is needed, as the rationality coefficient is otherwise not identifiable [22]. Similar findings are shown by [65], who model the rationality coefficient as a query-dependent function that might differ for the human labelers (see Section 5.1.6).
Another alternative to the rationality coefficient for representing irrational humans is achieved by introducing a query-independent error probability [58]. To be more precise, it is assumed that the human labeler only adheres to the Boltzmann distribution in (8) in 90% of cases and otherwise makes completely random decisions. This formulation is similar to Huber’s contaminated model [187].
@@ 5.1.3 Alternative Utility Notions
[143] show that the Boltzmann model does not generally lead to an identifiable reward function using (9) by presenting three concrete scenarios for which identification is not possible. The root cause of the non-identifiability is the usage of a trajectory’s return as the utility in (9). They, therefore, suggest using a trajectory’s regret as an alternative, which provably leads to identifiable rewards.
A trajectory’s regret is the negated sum of the optimal policy’s advantage over each state-action pair in the trajectory. Empirically, it has been shown that this modification improves the alignment of the learned strategy with human preferences. The downside of this alternative is that regret depends on the unknown optimal policy. Recently, it has also been suggested to consider $Q$-values of a human policy as the utilities [189], while [117] used differences of cost functions that depend on the available choice set and the human’s uncertainty.
@@ 5.1.4 Human Feedback Models Beyond Boltzmann
While the human feedback model based on the Boltzmann distribution is the most widely used model nowadays, other models have also been considered in the literature. In particular, for the probability in (4) other models such as the Thurstone model [269, 149, 28], the ridge-noise model [227], the binary model [245] or mixed forms thereof [273] have been considered. Of these models, only the Thurstone model [249] has a similar interpretation as the Boltzmann distribution based on perturbed utilities, only differing in the distribution of the perturbance random variables.
@@
[35] propose the Limiting Errors due to Similar Selections (LESS) model that is inspired by the attribute rule model suggested by [102]. It assumes a feature map for trajectories and a (similarity) function mapping trajectory features and trajectories to integers and uses a two-stage process for modeling the human feedback (or choice): First, choosing a trajectory feature according to the Boltzmann distribution and then a trajectory with the (logarithmic) similarity functions as the utilities within the Boltzmann distribution. Their experiments show that this model can capture human feedback more appropriately than the standard Boltzmann distribution.
@@
[2] evaluate the usage of a generative model for learning from human preferences. More specifically, instead of assuming some underlying utility as in the Bradley-Terry model, they attempt to train a model to generate the human feedback (inter-temporal preferences in this case, see Section 3.2.3) and directly interpret this feedback as reward. However, they found that this empirically does not perform as well as the inter-temporal Bradley-Terry model.
@@ 5.1.5 Misspecification
The human feedback model may be misspecified in various ways. [184] investigate the problem of misspecifying the nature of human feedback that can be either literal or pedagogical. The former means that the human gives targeted feedback for solving the actual RL problem, while the latter means that the human gives targeted feedback that is deemed helpful for the learner. They show theoretically and empirically that the case of a learner assuming a pedagogical with an actual literal human always performs worse than the reversed case, i.e., a learner assuming a literal with an actual pedagogical human.
A related question is studied by [77], namely, what if the learner makes incorrect assumptions about the choices from which the human selects its feedback? They consider different types of such choice set misspecification and show that depending on the type of misspecification, the performances might vary drastically, even leading to no losses at all in specific cases.
In the field of inverse RL, the general question of the robustness of reward learning in terms of a misspecified human feedback model is theoretically investigated by [239]. It turns out that the optimality model is not robust with respect to any misspecification, the Boltzmann model is robust for quite a range of specific misspecifications, and the degree of robustness of the maximal causal entropy model lies between the latter two. Even though these results are primarily derived for inverse RL, they also have similar immediate implications for RLHF.
@@ 5.1.6 Diverse Preferences
One potential issue with the RLHF framework is that it does not specify whose preferences to align to. In practice, it is common to aggregate preference feedback from multiple users and learn a single reward model to try to predict the mean preference. In contrast, [15] investigate this question more explicitly and proposes to learn multiple reward functions, which can then be aggregated in arbitrary manners and even be utilized to find consensus among people with different preferences. [188] consider the case of multiple humans that provide the feedback and use a mixture model of Plackett-Luce models to represent the feedback more accurately. With a stronger focus on the active retrieval of human feedback, [78] model the problem of selecting a suitable human labeler as a variant of the multi-armed bandit problem [152] that they call hidden-utility bandit. In this variant, the agent has in each decision round the choice between two things: (i) drawing a bandit arm and receiving a hidden arm-dependent utility and observing an item, or (ii) asking a human for a preference for two pairs of items but incurring a human-specific query cost. The feedback mechanism of an individual human is modeled via a Boltzmann distribution with a known individual rationality coefficient. The same modeling of human feedback is also considered by [16], who, however, use a Bayesian approach to determine which person should be queried in order to obtain the most informative feedback in expectation. [65] model the rationality coefficient as a query-dependent function that might differ for the human labelers.
@@ 5.1.7 Relaxation of the Markov Assumption
Most works assume that the human feedback is given based on a latent Markovian reward model, i.e., the return of a trajectory $\tau$ decomposes into a sum of independent rewards over state-action pairs (see (1)). [73] relax this assumption by dropping the need for the Markov property, such that the instantaneous reward might depend on hidden states. Similarly, [138] avoid the Markov assumption by utilizing a transformer as the preference model. A similar effect may be achieved by learning a state representation with a recurrent network in which the rewards are Markov, similar to the approach taken by [110], but we are not aware of any work exploring this. [2] work in a non-Markovian setting as well by utilizing memory-augmented networks for both the policy and the reward model.
@@ 5.2.1 Risk Minimization
The most prevalent variant for learning the reward function, already been presented in Section 2.3, is a special case of empirical risk minimization. The general approach of empirical risk minimization for reward function learning, assuming an underlying human feedback model with utilities $u_{l_{i}}$ as in (9), is to find the minimizer of
$\mathcal{L}(R;\mathcal{D})=\sum_{i=1}^{N}\ell\left(u_{l_{i}}(R),m(q_{i})\right)\,\text{,}$ (12)
@@
where $\mathcal{D}={\{(l_{i},q_{i})\}}_{i=1}^{N}$ is the given data set of observed label and query pairs and :×$\ell:\mathbb{R}\times\mathcal{Q}$ is a suitable loss function, where $\mathcal{Q}$ denotes the set of all possible label sets. To provide an example, let us consider the common case of pairwise trajectory comparisons. Here the grounding function $\psi$ is the identity, the queries are pairs of trajectories $q_{i}=\{\tau_{1}^{i},\tau_{2}^{i}\}$ and the labels $l_{i}\in\{\tau_{1}^{i}\succ\tau_{2}^{i},\tau_{1}^{i}\prec\tau_{2}^{i}\}=m(q_{i})$ are the human’s preference over the two trajectories. For a given query $q_{i}$, we then obtain (2) as a special case of (12) by using the loss function:
$\ell(u_{l_{i}}(R),m(q_{i}))$ $=-\log\left(\frac{1}{1+\exp(u_{l_{i}}(R)-u_{m(q_{i})\setminus l_{i}}(R))}\right)$
$=-\log\left(\frac{1}{1+\exp(\mathbb{E}_{\tau\sim\psi(l_{i})}[R(\tau)]-\mathbb{E}_{\tau\sim\psi(m(q_{i})\setminus l_{i})}[R(\tau)])}\right)\,\text{.}$
This is the negative log-likelihood for the Boltzmann distribution for the observational pair $(l_{i},m(q_{i}))$.
@@
For the entire learning process, a model class $\mathcal{R}$ is assumed for the reward function $R$. This model class is usually a parameterized class of functions, such as, for example, the class of linear reward functions [135]
=$\mathcal{R}=\{R_{\bm{\theta}}(s,a)=\bm{\theta}^{\top}\bm{\phi}(s,a)\,|\,(s,a)\in\mathcal{S}\times\mathcal{A},\bm{\theta}\in\mathbb{R}^{d}\},$
@@
where $\bm{\phi}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ is some state-action feature mapping. This entails that good features are known in advance such that rewards can be expressed as a linear combination of those features. Using such linear models may lead to reward model misspecification. Studying this setting, [33] propose to adapt the hyperparameter $\beta$ in Eq. 3 to account for this issue.
@@
Since the assumption of a linear reward model may be too strong in practice, most recent work is based on non-linear models, especially using differentiable models, but other cases have been investigated as well. In the former case, for instance, decision trees have been considered to learn an interpretable reward model (see Section 5.2.5). In the latter case, simple multilayer perceptron (MLP) has naturally been considered, but more recent deep learning architectures are more commonly used in the recent literature. Thus, especially with partially observable domains, the reward network may be composed of a state-action encoder followed by fully connected layers. For instance, [2] combine ResNet blocks for image processing, a learnable embedding table, a multi-modal transformer, LSTMs, and MLPs. Besides, [138] utilize a Transformer-based architecture [256], motivated by the observation that rewards are often non-Markovian.
@@
In addition to the usual empirical risk in (12), it is also typical, as in supervised machine learning, to add a regularization function to prevent overfitting:
$\mathcal{L}(R_{\bm{\theta}};\mathcal{D})=\sum_{i=1}^{N}\ell(u_{c_{i}}(R_{\bm{\theta}}),\mathcal{C}_{i})+\lambda_{r}(\bm{\theta})\,\text{,}$ (13)
@@
where $\lambda_{r}:\Theta\to\mathbb{R}_{+}$ is a regularization function defined on the parameter space $\Theta$ of the underlying reward model class. For instance, [58] simply use L2 regularization and also considers dropout in some domains.
@@
The main supervised loss to train the reward model can also be augmented with additional losses corresponding to auxiliary tasks to avoid overfitting and improve generalizability. For instance, [2] use a behavior cloning loss and add a policy head to the reward network, thereby preventing that the reward model drifts too far from its initialization from the policy.

### v2

@@ 5.1.1 Boltzmann Distribution
Human choice models as in (4) have been studied for a long time in various scientific fields such as psychology (Thurstone, 1927), economics (Train, 2009), or behavioral science (Cattelan, 2012). Accordingly, there are many different choice models to resort to for (5), which, in some cases, are the same models, just under different names. A popular class of such human choice models assumes every choice option $c$ to be equipped with a (latent) utility $u_{c}$, which the human perceives in a perturbed way. This perturbation is modeled by means of perturbation random variables $\epsilon_{c}$ that perturb the utility in an additive way, so that (4) becomes
@@
and we shall now stick to the RLHF translation from now on. These probabilities depend on the specific distributional assumptions that are made on the perturbation variables that only for specific cases lead to a closed-form of the right-hand sides of the latter equations. When stipulating a standard Gumbel distribution for the perturbations, one always obtains a closed form that is proportional to the exponential utility of the provided label:
@@
This is known as the Boltzmann distribution that also appears in a perhaps slightly modified version in various different subfields of ML and statistics. When restricting to discrete (choice) sets for $m(q)$, this distribution is also known as the multinomial logit model (Train, 2009) or Gibbs measure (Georgii, 2011), and as the Bradley-Terry model (Bradley & Terry, 1952) when the choice sets consist of pairs. All of these also have a link to the Plackett-Luce model (Luce, 1959; Plackett, 1975), which is a probability distribution on the space of total orders or rankings (see Alvo & Yu (2014) for details).
@@
This model is often used for various reasons. A particularly compelling reason is the closed analytic form, which in turn makes it possible to obtain a closed form for the gradient with respect to the utilities. Another reason is that this model satisfies Luce’s axiom of choice (Plackett, 1975), which requires the probability of choosing an option from a pool of choice options not being affected by the presence or absence of other options in the pool. In this way, coherent decision-making is ensured, which, however, might be challenged as humans are likely not making fully rational decisions (see Section 4.2.1). Jeon et al. (2020) show that the usage of the Boltzmann distribution is justified by the principle of maximum entropy. More precisely, they show that it is the maximum entropy distribution over choices for a so-called satisficing human decision maker, i.e., one who is making in expectation a choice with an optimal reward up to some slack $\epsilon>0$.
@@
To build the bridge between reward learning and the modeling of human feedback, the Boltzmann distribution can be used by assuming that the utilities can be represented as a function of the reward function, usually featuring the return of a trajectory. More specifically, one assumes a grounding function $G$ that maps choice options (or labels) to the set of distributions over trajectories and sets the utility of a label $l$ as
$u_{l}:=\mathbb{E}_{\tau\sim G(l)}[R(\tau)]\,\text{.}$ (9)
Note that $u_{l}$ depends on the return $R,$ so that we also may use $u_{l}(R)$ to emphasize this dependency. For the common case of pairwise trajectory comparisons, where for two trajectories $\tau_{1},\tau_{2}$ we obtain for the possible labels $l\in\{\tau_{1}\succ\tau_{2},\tau_{1}\prec\tau_{2}\}$ the respective utility by using the projection onto the preferred trajectory as the grounding function $G.$ Accordingly, the utility of the label represents essentially the utility of the preferred trajectory of that label, i.e., $\tau_{1}$ or $\tau_{2}$ in this case. As another example consider the case of e-stops feedback (see Section 3.2.2). Here, the possible labels $l$ provided by the user are $\texttt{STOP}_{t}$ and CONT encoding the stopping at time $t$ or the continuation of a trajectory. For the grounding function, one can define
$G(l)=\begin{cases}\tau,&l=\texttt{CONT},\\ \tau_{0:t}\tau_{t}\dots\tau_{t},&l=\texttt{STOP}_{t},\end{cases}$
@@
where $\tau_{0:t}$ is the trajectory of $\tau$ trimmed to the stopping time $t$ and $\tau_{t}$ is the action-state pair at time $t.$ Table 1 in Jeon et al. (2020) provides an overview of the different grounding functions that lead to a specific feedback type. It is worth noting that one can also easily find a grounding function for the feedback type of a (partial) order over trajectories as considered, for instance, by Myers et al. (2021). Moreover, one can generalize this modeling approach by using (partial) segments instead of trajectories.
@@
Although this general human feedback model has been much in use and shown to be useful for the sake of human alignment, it is not without its critics (see Lindner & El-Assady (2022) or Section 3.2.1 in Casper et al. (2023)). This has led to different adaptions of the general model based on the Boltzmann distribution that will be discussed in the following. Moreover, we will also concisely review other human feedback models that have been in use besides the Boltzmann distribution, discuss relevant work on the consequences or robustness of human feedback model misspecification, and highlight contributions on varying the standard assumptions on the nature of the human feedback.
@@ 5.1.2 Human-Specific Rationality Coefficient
The higher $\beta$, the more (10) resembles a pointmass distribution modeling a highly rational human labeler (decision-maker) that is always able to identify the option the with highest utility, while the lower $\beta$, the more (10) resembles a uniform distribution modeling a highly irrational human labeler (decision-maker) acting purely at random. Without this extension, the commonly considered Boltzmann distribution (or Bradley-Terry model in the common case of pairwise comparisons) in (8) assumes a rationality coefficient of $1$. Ghosal et al. (2023) show in their experiments that the estimation of this coefficient can indeed positively influence reward learning. For the estimation, however, a calibration reward function is needed, as the rationality coefficient is otherwise not identifiable (Bengs & Hüllermeier, 2020). Similar findings are shown by Daniels-Koch & Freedman (2022), who model the rationality coefficient as a query-dependent function that might differ for the human labelers (see Section 5.1.6).
Another alternative to the rationality coefficient for representing irrational humans is achieved by introducing a query-independent error probability (Christiano et al., 2017). To be more precise, it is assumed that the human labeler only adheres to the Boltzmann distribution in (8) in 90% of cases and otherwise makes completely random decisions. This formulation is similar to Huber’s contaminated model (Mu & Xiong, 2023).
@@ 5.1.3 Alternative Utility Notions
Knox et al. (2024) show that the Boltzmann model does not generally lead to an identifiable reward function using (9) by presenting three concrete scenarios for which identification is not possible. The root cause of the non-identifiability is the usage of a trajectory’s return as the utility in (9). They, therefore, suggest using a trajectory’s regret as an alternative, which provably leads to identifiable rewards.
A trajectory’s regret is the negated sum of the optimal policy’s advantage over each state-action pair in the trajectory. Empirically, it has been shown that this modification improves the alignment of the learned strategy with human preferences. The downside of this alternative is that regret depends on the unknown optimal policy. Recently, it has also been suggested to consider $Q$-values of a human policy as the utilities (Myers et al., 2023), while Holladay et al. (2016) used differences of cost functions that depend on the available choice set and the human’s uncertainty.
@@ 5.1.4 Human Feedback Models Beyond Boltzmann
While the human feedback model based on the Boltzmann distribution is the most widely used model nowadays, other models have also been considered in the literature. In particular, for the probability in (4) other models such as the Thurstone model (Wilson et al., 2012; Kupcsik et al., 2018; Bıyık et al., 2020), the ridge-noise model (Schoenauer et al., 2014), the binary model (Sugiyama et al., 2012) or mixed forms thereof (Wirth et al., 2016) have been considered. Of these models, only the Thurstone model (Thurstone, 1927) has a similar interpretation as the Boltzmann distribution based on perturbed utilities, only differing in the distribution of the perturbance random variables.
@@
Bobu et al. (2020b) propose the Limiting Errors due to Similar Selections (LESS) model that is inspired by the attribute rule model suggested by Gul et al. (2014). It assumes a feature map for trajectories and a (similarity) function mapping trajectory features and trajectories to integers and uses a two-stage process for modeling the human feedback (or choice): First, choosing a trajectory feature according to the Boltzmann distribution and then a trajectory with the (logarithmic) similarity functions as the utilities within the Boltzmann distribution. Their experiments show that this model can capture human feedback more appropriately than the standard Boltzmann distribution.
@@
Abramson et al. (2022) evaluate the usage of a generative model for learning from human preferences. More specifically, instead of assuming some underlying utility as in the Bradley-Terry model, they attempt to train a model to generate the human feedback (inter-temporal preferences in this case, see Section 3.2.1) and directly interpret this feedback as reward. However, they found that this empirically does not perform as well as the inter-temporal Bradley-Terry model.
@@ 5.1.5 Misspecification
The human feedback model may be misspecified in various ways. Milli & Dragan (2020) investigate the problem of misspecifying the nature of human feedback that can be either literal or pedagogical. The former means that the human gives targeted feedback for solving the actual RL problem, while the latter means that the human gives targeted feedback that is deemed helpful for the learner. They show theoretically and empirically that the case of a learner assuming a pedagogical feedback with an actual literal human always performs worse than the reversed case, i.e., a learner assuming a literal feedback with an actual pedagogical human.
A related question is studied by Freedman et al. (2021), namely, what if the learner makes incorrect assumptions about the choices from which the human selects its feedback? They consider different types of such choice set misspecification and show that depending on the type of misspecification, the performances might vary drastically, even leading to no losses at all in some specific cases.
In the field of inverse RL, the general question of the robustness of reward learning in terms of a misspecified human feedback model is theoretically investigated by Skalse & Abate (2023). It turns out that the optimality model is not robust with respect to any misspecification, the Boltzmann model is robust for quite a range of specific misspecifications, and the degree of robustness of the maximal causal entropy model lies between the latter two. Even though these results are primarily derived for inverse RL, they also have similar immediate implications for RLHF.
@@ 5.1.6 Diverse Preferences
One potential issue with the RLHF framework is that it does not specify whose preferences to align to. It is common to request feedback from multiple labelers in a crowd-sourcing setting, in which case the different labelers may disagree. There are two main ways to deal with this challenge: Either trying to learn each labeler’s preference separately, or trying to learn a model of the group’s mean preference.
Bakker et al. (2022) investigate the first option by proposing to learn multiple reward functions, which can then be aggregated in arbitrary manners and even be utilized to find consensus among people with different preferences. The second is more commonly used, however: Xue et al. (2023a) learn a single reward function from multiple humans who may give diverse and inconsistent feedback, aiming to stabilize learning in spite of these inconsistencies using regularization, a consistency constraint, and ensembling. Similarly, Chhan et al. (2024) try to estimate the correct preference for pairwise trajectories directly by combining the users’ expressed preference labels instead of assuming individual reward functions. As a middle-ground between the single reward function and multiple ones, Myers et al. (2021) propose to learn a multimodal reward function that captures multiple people’s preferences and use a mixture model of Plackett-Luce models to represent the feedback more accurately.
With a stronger focus on the active retrieval of human feedback, Freedman et al. (2023) model the problem of selecting a suitable human labeler as a variant of the multi-armed bandit problem (Lattimore & Szepesvári, 2020), which they call hidden-utility bandit. In this variant, the agent has in each decision round the choice between two options: (i) drawing a bandit arm, then receiving a hidden arm-dependent utility, and finally observing an item, or (ii) querying a human to observe a preference between two sampled items and incurring a human-specific query cost. The feedback mechanism of all human teachers is modeled via a same Boltzmann distribution, differing only in their known individual rationality coefficients. The same modeling of human feedback is also considered by Barnett et al. (2023), who, however, use a Bayesian approach to determine which person should be queried in order to obtain the most informative feedback in expectation. Daniels-Koch & Freedman (2022) investigate the rationality coefficient already considered in the previously mentioned work and model it as a query-dependent function that might differ for the human labelers.
@@ 5.1.7 Relaxation of the Markov Assumption
Most works assume that the human feedback is given based on a latent Markovian reward model, i.e., the return of a trajectory $\tau$ decomposes into a sum of independent rewards over state-action pairs (see (1)). Early et al. (2022) relax this assumption by dropping the need for the Markov property, such that the instantaneous reward might depend on hidden states. Similarly, Kim et al. (2023) avoid the Markov assumption by utilizing a transformer as the preference model. A similar effect may be achieved by learning a state representation with a recurrent network in which the rewards are Markov, similar to the approach taken by Hafner et al. (2023), but we are not aware of any work exploring this. Abramson et al. (2022) work in a non-Markovian setting as well by utilizing memory-augmented networks for both the policy and the reward model.
@@ 5.2.1 Empirical Risk Minimization
The most prevalent variant for learning the reward function, already been presented in Section 2.3, is a special case of empirical risk minimization. The general approach of empirical risk minimization for reward function learning, assuming an underlying human feedback model with utilities as in (9) is to find the minimizer of
$\mathcal{L}(R;\mathcal{D})=\sum_{i=1}^{N}\ell\left(u_{\cdot}(R),l_{i},m(q_{i})\right)\,\text{,}$ (12)
@@
where $\mathcal{D}={\{(l_{i},q_{i})\}}_{i=1}^{N}$ is the given data set of observed label and query pairs, :×$\ell:\mathbb{R}\times\mathcal{Q}\times\mathcal{C}$ is a suitable loss function with $\mathcal{Q}$ being the set of all possible label sets, and $u_{\cdot}(R)$ denoting the utility (depending on the return $R$) of the possible labels for the given label-query pair $l_{i},m(q_{i}).$ As an illustration, consider the common case of pairwise trajectory comparisons where queries are pairs of trajectories $q_{i}=\{\tau_{1}^{i},\tau_{2}^{i}\}$, and labels $l_{i}\in\{\tau_{1}^{i}\succ\tau_{2}^{i},\tau_{1}^{i}\prec\tau_{2}^{i}\}=m(q_{i})$ are the human’s preference over the two trajectories. For a given query $q_{i}$, we then obtain (2) as a special case of (12) by using the loss function:
$\ell(u_{\cdot}(R),l_{i},m(q_{i}))$ $=-\log\left(\frac{1}{1+\exp(u_{m(q_{i})\setminus l_{i}}(R)-u_{l_{i}}(R))}\right)$
$=-\log\left(\frac{1}{1+\exp(\mathbb{E}_{\tau\sim G(m(q_{i})\setminus l_{i})}[R(\tau)]-\mathbb{E}_{\tau\sim G(l_{i})}[R(\tau)])}\right)\,\text{,}$
where in the case of pairwise comparison, $u_{\cdot}(R)=(u_{l_{i}}(R),u_{m(q_{i})\setminus l_{i}}(R))$ and the grounding function $G$ is the projection onto the preferred trajectory. This is the negative log-likelihood for the Boltzmann distribution for the observational pair $(l_{i},m(q_{i}))$.
@@
For the entire learning process, a model class $\mathcal{R}$ is assumed for the reward function $R$. This model class is usually a parameterized class of functions, such as, for example, the class of linear reward functions (Katz et al., 2021)
=$\mathcal{R}=\{R_{\bm{\psi}}(s,a)=\bm{\psi}^{\top}\bm{\phi}(s,a)\,|\,(s,a)\in\mathcal{S}\times\mathcal{A},\bm{\psi}\in\mathbb{R}^{d}\},$
@@
where $\bm{\phi}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ is some state-action feature mapping. This entails that good features are known in advance such that rewards can be expressed as a linear combination of those features. Using such linear models may lead to reward model misspecification. Studying this setting, Bobu et al. (2020a) propose to adapt the hyperparameter $\beta$ in (3) to account for this issue.
@@
Since the assumption of a linear reward model may be too strong in practice, most recent work is based on non-linear models, especially using differentiable models, but other cases have been investigated as well. In the latter case, for instance, decision trees have been considered to learn an interpretable reward model (see Section 5.2.4). In the former case, simple multilayer perceptron (MLP) has naturally been considered, but more recent deep learning architectures are more commonly used in the recent literature. Thus, especially with partially observable domains, the reward network may be composed of a state-action encoder followed by fully connected layers. For instance, Abramson et al. (2022) combine ResNet blocks for image processing, a learnable embedding table, a multi-modal transformer, LSTMs, and MLPs. Besides, Kim et al. (2023) utilize a Transformer-based architecture (Vaswani et al., 2017), motivated by the observation that rewards are often non-Markovian.
@@
In addition to the usual empirical risk in (12), it is also typical, as in supervised ML, to add a regularization function to prevent overfitting:
$\mathcal{L}(R_{\bm{\psi}};\mathcal{D})=\sum_{i=1}^{N}\ell(u_{l_{i}}(R_{\bm{\psi}}),l_{i},m(q_{i}))+\lambda_{r}(\bm{\psi})\,\text{,}$ (13)
@@
where $\lambda_{r}:\Psi\to\mathbb{R}_{+}$ is a regularization function defined on the parameter space $\Psi$ of the underlying reward model class. For instance, Christiano et al. (2017) simply use L2 regularization and also consider dropout in some domains. Recently, Verma & Metcalf (2024) propose to define a more complex regularization term, which consists in biasing the learned rewards to be proportional to an approximate state importance provided by a trained Transformer-based forward model.
@@
The main supervised loss to train the reward model can also be augmented with additional losses corresponding to auxiliary tasks to avoid overfitting and improve generalizability. For instance, Abramson et al. (2022) use a behavior cloning loss and add a policy head to the reward network, thereby preventing that the reward model drifts too far from its initialization from the policy. Metcalf et al. (2023) design a reward model using state-action representations trained to be temporally consistent via self-supervised learning. On a related note, the scalar preference optimization problem has been extended to a multidimensional one by Zhong et al. (2024) to represent diverse human preferences and Marta et al. (2023) for query efficiency.